OpenStudy (anonymous):

Can someone please help me how to solve this what is the y coordinate of the y intercept line that passes through the points (-3,-1) and (1,7) Do you have to do y2-y1/x2-x1?

OpenStudy (anonymous):

Yes you need the slope. Then you need to create the formula in standard form

OpenStudy (anonymous):

@nelsonjedi after I find the slope what do I do next?

OpenStudy (anonymous):

Because I got 2 for the slope

OpenStudy (anonymous):

Use either point to create the formula for your line.

OpenStudy (anonymous):

Is it y=2x+?

OpenStudy (anonymous):

Oh okay thanks

OpenStudy (anonymous):

Is it -7

OpenStudy (anonymous):

(y-y1)= m(x-x1)

OpenStudy (anonymous):

Yes I see is that the final answer or am I supposed to add 7 to both sides?

OpenStudy (anonymous):

Sorry I goofed m= slope = 2. So we have (y-y1)=2(x-x1)..

OpenStudy (anonymous):

Now plug in either point I will use (1,7) so we have (y-7)=2(x-1) or y-7=2x-2.

OpenStudy (anonymous):

Now add 7 to both sides and you get y= 2x+5

OpenStudy (anonymous):

I got y=2x+2 is that right?

OpenStudy (anonymous):

To find the x intercept you make y=0 and solve for x. To find the y intercept you make x =0 and solve for y

OpenStudy (anonymous):

7+(-2)= 5 y=2x+5

OpenStudy (anonymous):

Where'd you get the -2?

OpenStudy (anonymous):

On the right side of the equation we have 2(x-1) we have to distribute the 2 so we end up with 2x-2. Now we have y-7=2x-2. Now add 7 to both sides

OpenStudy (anonymous):

Oh okay that makes sense thanks so much!!!